Secret:

Find all points, if any, where y - 4x = 12 intersects 2 - y = 2(x + 2)2 There are no points of intersection (-3, 0) (3, 0) (-3, 0) and (3, 0)

Shadow:

@Secret How far have you gotten on this one?

Secret:

Not much, i need help, i'm not to sure on how to do this...

Shadow:

Do you have graphing paper?

Shadow:

\[y - 4x = 12\] \[2 - y = 2(x + 2)^2\] This is what they look like, correct?

Shadow:

\[y = 4x + 12\] \[y = -2(x +2)^2 + 2\]

Shadow:

You can throw the points in https://www.desmos.com/calculator and graph them.

Shadow:

Throw the graphs*

bm717:

Throw the equations*